# Fair Concurrent Training of Multiple Models in Federated Learning

### v1

@@ Abstract
Federated learning (FL) enables collaborative learning across multiple clients. In most FL work, all clients train a single learning task. However, the recent proliferation of FL applications may increasingly require multiple FL tasks to be trained simultaneously, sharing clients’ computing and communication resources, which we call Multiple-Model Federated Learning (MMFL). Current MMFL algorithms use naïve average-based client-task allocation schemes that can lead to unfair performance when FL tasks have heterogeneous difficulty levels, e.g., tasks with larger models may need more rounds and data to train. Just as naïvely allocating resources to generic computing jobs with heterogeneous resource needs can lead to unfair outcomes, naïve allocation of clients to FL tasks can lead to unfairness, with some tasks having excessively long training times, or lower converged accuracies. Furthermore, in the FL setting, since clients are typically not paid for their training effort, we face a further challenge that some clients may not even be willing to train some tasks, e.g., due to high computational costs, which may exacerbate unfairness in training outcomes across tasks. We address both challenges by firstly designing FedFairMMFL, a difficulty-aware algorithm that dynamically allocates clients to tasks in each training round. We provide guarantees on fairness and FedFairMMFL’s convergence rate. We then propose a novel auction design that incentivizes clients to train multiple tasks, so as to fairly distribute clients’ training efforts across the tasks. We show how our fairness-based learning and incentive mechanisms impact training convergence and finally evaluate our algorithm with multiple sets of learning tasks on real world datasets.
@@ I Introduction
Federated learning (FL) enables multiple devices or users (called clients) to collectively train a machine learning model  [1, 2, 3, 4], with data staying at the user side. Several recent works have demonstrated FL’s potential in many applications, such as smart homes [2, 3], analyzing healthcare data [5], and next word prediction [6]. Despite the large and growing set of FL applications, however, the vast majority of FL research assumes that each client only trains one learning task during FL. In practice, clients may concurrently train multiple tasks, e.g., if smartphones contribute to training both next-word prediction and advertisement recommendation models, or if hospitals contribute to training separate prediction models for heart disease and breast cancer. Clients could also train multiple versions of a model, in order to select the best-performing one for deployment, e.g., training many different neural network architectures on the same data. We call this setting in which FL clients simultaneously train multiple tasks, Multiple-Model Federated Learning (MMFL) [7, 8, 9, 10, 11]. Training multiple tasks in one FL setting amortizes the real-world costs of recruiting clients and coordinating the training time across clients, improving resource efficiency.
@@
Since FL clients, especially mobile or Internet of Things devices, may be resource-constrained [1], they may not be able to train every FL task at the same time. Thus, we must allocate clients to the different tasks in each training round. Prior work on MMFL has shown that with careful selection of clients for each task in each training round, concurrent training yields comparable accuracy to training tasks separately [7]. These works, however, seek to optimize the average task accuracy or training time, and often assume that training tasks have similar difficulties, e.g., training multiple copies of the same model [8, 7]. In practice, however, FL tasks may be requested by different entities and users, and it is important to ensure comparable convergence time and converged accuracies across tasks of varied difficulty levels. Hospitals, for example, may wish to concurrently train prediction models for different diseases, without favoring one disease’s model over another. Likewise, in applications like autonomous driving, improving the worst-performing FL task helps to improve the overall quality of experience for users. Indeed, even if the multiple tasks are different models trained by the same FL operator, the operator will not know which model performs best until the training of all models is complete: we therefore would want to dynamically allocate clients to tasks so as to optimize the performance of the worst-performing task, which is accomplished by optimizing for fairness. To the best of our knowledge, we are the first to extend the MMFL setting to consider fair training of concurrent FL models.
@@ I-A Research Challenges in Fair MMFL
In considering fairness across multiple different models, we seek to ensure that each model achieves comparable training performance, in terms of their converged accuracies, or the time taken to reach threshold accuracy levels. The first challenge to doing so is that the multiple models can have varying difficulties and thus require different amounts of data and computing resources to achieve the same accuracy. Existing MMFL training algorithms, e.g., the round-robin algorithm proposed in [7], allocate clients to tasks uniformly, or assume that tasks have similar or the same difficulty, e.g., training multiple copies of the same model [8]. However, while such strategies arguably achieve fairness, or at least equality, in resource (i.e., client) allocation to different models, they may lead to “unfair” outcomes. For example, smart home devices may train simple temperature prediction models that take low-dimensional timeseries as input data, as well as complex person identification ones that use high-dimensional image input data [3]. Allocating equal resources to each model may then yield poor performance for the more difficult ones.
Much work in computer systems and networks has focused on fairness in allocating resources to different computing jobs or network flows, given their heterogeneous resource requirements [12, 13, 14]. Simply applying these allocation algorithms to MMFL, however, is difficult: here, each “resource” is a client, and the performance improvement realized by allocating this client to a given model’s training task is unknown in advance and will change over time. This is because it depends on the current model state and the data distribution across clients, both of which are highly stochastic. We must therefore handle significant uncertainty and dynamics in how much each client contributes to each task’s training performance.
Finally, purely viewing MMFL fairness as a matter of assigning clients to tasks ignores the fact that clients may not be equally willing to contribute resources to each model. For example, smartphone users may benefit from better next-word prediction models if they text frequently, but may derive limited benefit from a better ad recommendation model. Users may refuse to compute updates if they judge that associated computing costs outweigh personal benefits. If multiple users prefer one model over another, then the disfavored model will overfit to users willing to train it, again leading to an unfair outcome. In an extreme case, no users may be assigned to train this task at all. Prior work[15] has proposed mechanisms to incentivize clients to help train multiple non-simultaneously trained FL models, but the focus was on optimizing the task take-up rate, instead of ensuring that clients are fairly distributed across models.
@@ I-B Our Contributions
Our goal is to design methods to ensure fair concurrent training of MMFL, addressing the challenges above. To do so, we utilize the popular metric of $\alpha$-fairness, commonly used for fair resource allocation in networking [16] and previously proposed to measure fairness across clients in training a single FL model [17]. We use $\alpha$-fairness and Max-Min fairness (a special case of $\alpha$-fairness when $\alpha\rightarrow\infty$ [16]) to design novel algorithms that both incentivize clients to contribute to FL tasks and adaptively allocate these incentivized clients to multiple tasks during training. Intuitively, these algorithms ensure that client effort is spread across the MMFL tasks in accordance with their difficulty, helping to achieve fairness with respect to task performances. After discussing related works in Section II, we make the following contributions:
  • We first suppose that all clients are willing to train all models, and propose FedFairMMFL, an algorithm to fairly allocate clients to tasks (i.e., models) in Section III. Our algorithm dynamically allocates clients to tasks according to the current performance levels of all tasks. We design FedFairMMFL to approximately optimize the $\alpha$-fairness of all tasks’ training accuracies.
  • From the perspective of a single model, our client-task allocation algorithm in MMFL can be viewed as a form of client selection. We therefore adapt and refine convergence guarantees from analyses of client selection in FL to analyze the convergence of MMFL tasks under FedFairMMFL in Section IV, showing that FedFairMMFL preferentially accelerates convergence of more difficult tasks (Theorem 4 and Corollary 5) and allows each model’s training to converge (Corollary 6).
@@
  • We finally validate our client incentivization and allocation algorithms on multiple FL datasets. Our results show that our algorithm achieves a higher minimum accuracy across tasks than random allocation and round robin baselines, and a lower variance across task accuracy levels, while maintaining the average task accuracy. Our max-min client incentivization further improves the minimum accuracy in the budget-constrained regime.
@@ II Related Work
We distinguish our MMFL setting from that of multi-task learning (MTL), in which multiple training tasks have the same or a similar structure [18], e.g., a similar underlying lower rank structure. Unlike clustered FL [19, 20], our work is different from combining MTL with FL, instead we consider multiple distinct learning tasks. Our analysis of MMFL performance is closer to prior convergence analysis of client selection in FL [21, 22, 23]. We adapt prior work on FL convergence to reveal the relationships between model convergence and our definition of fairness across tasks.
Past work on multi-model federated learning also aims to optimize the allocation of clients to model training tasks. To the best of our knowledge, we are the first to explicitly consider fairness across the different tasks. Some algorithms focus on models with similar difficulties, for which fairness is easier to ensure [7], while other works propose reinforcement learning [24], heuristic [9], or asynchronous client selection algorithms [10] to optimize average model performance. Other work optimizes the training hyperparameters [25] and allocation of communication channels to tasks [21].
@@ III Fair Client Task Allocation for Multiple-Model FL
We aim to train multiple models (i.e., tasks) [^1] We use model and task interchangeably in this paper. concurrently, such that there is fair fair training performance across models, e.g., in terms of the time taken, or converged accuracies. Models may be trained on distinct or the same datasets. This performance is a function of which clients train which model at which training round. In this section, we assume clients always train any given model upon request and propose a client-model (i.e., client-task) allocation scheme that enables fair training of the multiple tasks. In Sec. V, we introduce auction mechanisms, which incentivize clients to commit to tasks at a fair rate across tasks, before training begins (Fig. 1).
@@
MMFL framework: Consider $K$ clients collaboratively training $S$ distinct machine learning tasks in MMFL (Fig. 1). If $S=1$, our setting reduces to traditional FL. Each client is equipped with local data for training each task. All clients can communicate with a central server that maintains a global model $\textbf{w}_{s}$ for each task $s\in\mathbb{S}$. Let each client $k$’s local loss over its local dataset $\mathcal{D}_{k,s}$ for task $s$ be $F_{k,s}(\textbf{w}_{s})=\sum_{d\in\mathcal{D}_{k,s}}\ell\left(d,\textbf{w}_{s}\right)/\left|\mathcal{D}_{k,s}\right|$, where $\ell$ is a loss function that measures the accuracy of the model $\textbf{w}_{s}$ on data point $d$. We then define the global loss for task $s$ as a weighted sum over the local losses: $f_{s}(\textbf{w}_{s})=\sum_{k=1}^{K}p_{k,s}F_{k,s}(\textbf{w}_{s})$, where $p_{k,s}=\frac{\left|\mathcal{D}_{k,s}\right|}{\sum_{k}\left|\mathcal{D}_{k,s}\right|}$ is the fraction of training data for task $s$ that resides at client $k$. Similar to traditional FL, in MMFL, the usual aim is to minimize the sum of losses [10]:
@@
Fairness over tasks: We introduce fairness into this objective by utilizing the celebrated $\alpha$-fair utility function often employed to ensure fair network resource allocation [16, 13]. Viewing the loss of each model as an inverse “utility,” we then define our $\alpha$-fair MMFL objective as
$\min_{\textbf{w}_{1},\textbf{w}_{2},...\textbf{w}_{S}}g^{\alpha}\left(\textbf{w}_{1},\textbf{w}_{2},...\textbf{w}_{S}\right)\equiv\sum_{t=1}^{S}f_{s}^{\alpha}(\textbf{w}_{t}),$ (2)
@@
To solve (2) within the FL framework, we will use gradient descent to design a training algorithm given the gradient
$\nabla_{\textbf{w}_{s}}g=\alpha f_{s}^{\alpha-1}(\textbf{w}_{s})\sum_{k}p_{k}\nabla F_{k,s}(\textbf{w}_{s}).$ (3)
FedFairMMFL training (Algorithm 1): Our MMFL training begins with the initialization of the global model weights $\textbf{w}_{s}$ for all tasks, along with an initial random client-task allocation (line 1). The algorithm then enters a loop over global rounds (lines 2-15). During each global round $t$, $m=C\times|\mathcal{C}|$ clients are active (line 3). Here $C$ is the rate at which clients are active. For all active clients, task allocation happens. We let $Sel_{s}(t)$ denote the set of clients allocated to task $s$ in round $t$, dropping the subscript $s$ if it is clear from context. Following prior work on MMFL [7, 8, 24], we allocate each client to at most one task during each training round, i.e., the $Sel_{s}(t)$ are disjoint across tasks $s$. For example, mobile clients or other resource constrained devices may be performing other background tasks, e.g., running multiple applications, and thus may not have the computing power or battery to complete training rounds of multiple tasks.
Intuitively, uniform client-task allocation methods, like random or round robin allocation [8], will train simpler tasks faster, since they allocate an equal number of clients to each task, in expectation. We would instead like to ensure tasks are trained at more similar rates (i.e., achieve more fair performance) with a difficulty-aware client-task allocation algorithm, in which more users are allocated to tasks that currently have a higher loss compared to other tasks. We can do so by inspecting the gradient of our MMFL objective in (3). Since resource-constrained clients can only compute gradients for one task per training round, we cannot exactly compute these gradients. Instead, we will assign each client to contribute to the gradient $\nabla F_{k,s}(\textbf{w}_{s})$ for each task $s$ at a probability in proportion to its weight for that task, i.e., proportional to $\alpha f_{s}(\textbf{w}_{s})^{\alpha-1}p_{k}$. We therefore allocate a task to each client according to the probability vector (lines 4-5):
@@
This allocation fulfills our intuition: tasks currently with higher losses $f_{s}$ have a higher probability of being allocated clients. For example, when $\alpha\rightarrow\infty$, all clients will be allocated to the worst performing task, and when $\alpha=1$, allocation will be uniform. In Section IV, we formalize this intuition by deriving a convergence bound for each task using FedFairMMFL. This selection scheme is unbiased across clients: each client is equally likely to train a given task.
@@
Given the clients allocated to each task $s$ at time $t$, $Sel_{s}(t)$, the server then sends the latest global model weights $\textbf{w}_{s}(t)$ to task $s$’s allocated clients (lines 7-8). Each client then trains its task on its corresponding local dataset by running $\tau$ SGD steps on the local loss $F_{k,s}$ of its assigned task $s$, as in traditional FL [1] (line 9). Let $\textbf{w}_{k,s}(t)$ denote the updated local model for active client $k$ on its allocated task $s$. Each active client then uploads $\textbf{w}_{k,s}(t)$ to the server (line 9), which aggregates the model parameters to synchronize the global model. This aggregation uses weights $p_{k,Sel}=p_{k}/\sum_{k\in Sel}p_{k^{\prime}}$ for each client $k$ (line 12), where $Sel$ is the set of clients training task $s$ in this training round. The server then obtains the global loss for each task $f_{s}$, to guide allocation in the next training round (line 13). The next training round then begins.
@@
2:for $Globalepoch=1,2,\ldots$ do
3:    $m=C\times|\mathcal{C}|$ clients are active.
@@
7:         Server sends latest global model weights of task $s$,
8:         $\textbf{w}_{s}$ to clients allocated.
9:         Client performs local SGD (stochastic gradient descent) $\tau$ times for allocated task, optimizing $\textbf{w}_{k,s}$, and uploads $\textbf{w}_{k,s}$ to the server.
10:    end for
11:    for all tasks $s\in\mathbb{S}$ do
12:         Server aggregates $\textbf{w}_{s}\leftarrow\textbf{w}_{s}-\eta_{t}\sum_{k}p_{k,Sel}\nabla F_{k}(\textbf{w}_{s})$.
13:         Obtain global accuracy $f_{s}$ for task $s$.
14:    end for

### v2

@@ Abstract
Federated learning (FL) enables collaborative learning across multiple clients. In most FL work, all clients train a single learning task. However, the recent proliferation of FL applications may increasingly require multiple FL tasks to be trained simultaneously, sharing clients’ computing resources, which we call Multiple-Model Federated Learning (MMFL). Current MMFL algorithms use naïve average-based client-task allocation schemes that often lead to unfair performance when FL tasks have heterogeneous difficulty levels, as the more difficult tasks may need more client participation to train effectively. Furthermore, in the MMFL setting, we face a further challenge that some clients may prefer training specific tasks to others, and may not even be willing to train other tasks, e.g., due to high computational costs, which may exacerbate unfairness in training outcomes across tasks. We address both challenges by firstly designing FedFairMMFL, a difficulty-aware algorithm that dynamically allocates clients to tasks in each training round, based on the tasks’ current performance levels. We provide guarantees on the resulting task fairness and FedFairMMFL’s convergence rate. We then propose novel auction designs that incentivizes clients to train multiple tasks, so as to fairly distribute clients’ training efforts across the tasks, and extend our convergence guarantees to this setting. We finally evaluate our algorithm with multiple sets of learning tasks on real world datasets, showing that our algorithm improves fairness by improving the final model accuracy and convergence speed of the worst performing tasks, while maintaining the average accuracy across tasks.
@@ I Introduction
Federated learning (FL) enables multiple devices or users (called clients) to collectively train a machine learning model  [1, 2, 3, 4, 5], with data staying at the user side. Several recent works have demonstrated FL’s potential in many applications, such as smart homes [2, 3], analyzing healthcare data [6], and next word prediction [7]. Despite the expanding range of FL applications, however, the vast majority of FL research assumes that clients jointly train just one model during FL [8, 9, 10]. In practice, clients may increasingly need to train multiple tasks (models) at the same time, e.g., smartphones contributing to training both next-word prediction and advertisement recommendation models; data from wearables contributing to the FL training of various human physical activity recognition and health monitoring (e.g. heart rate prediction or early disease detection) models concurrently; hospitals contributing to training separate prediction models for heart disease and breast cancer; or intelligent vehicles generating data for driving-related tasks (e.g. control and localization), and safety-related tasks (e.g., obstacle detection and driver monitoring). Clients could also train multiple versions of a model, in order to select the best-performing one for deployment, e.g., training many different neural network architectures on the same data. While a naïve solution would simply train all models sequentially, doing so is inefficient, as the required training time would scale linearly in the number of models [11]. This approach also makes the later models wait for the prior models to finish training, which may be unfair. Moreover, training multiple tasks concurrently in one FL setting amortizes the real-world costs of recruiting clients and coordinating the training time across clients, improving resource efficiency. We call this setting in which FL clients concurrently train multiple tasks, Multiple-Model Federated Learning (MMFL) [12, 13, 14, 15, 16, 17, 11], and some prior research has demonstrated that this concurrent training approach can accelerate the overall training process[12, 13].
Prior works on MMFL [13, 12, 15, 14, 16] seek to optimize the client-task allocation each training round, given the resource constraints of devices. These works allocate clients to tasks so as to optimize the average task accuracy or training time, and often assume that training tasks have similar difficulties, e.g., training multiple copies of the same model [13, 12].
@@
In practice, however, FL tasks may have varying difficulty levels, and may be requested by different entities and users. Therefore, it is important to ensure comparable convergence time and converged accuracies across tasks of varied difficulty levels. For instance, wearable devices frequently train lightweight models for monitoring simpler metrics like step counting or heart rate tracking, concurrently with more complex tasks like early disease detection (e.g., arrhythmia detection). We wish to concurrently train these different models, without favouring one model over the other. Likewise, in applications like autonomous driving, improving the worst-performing FL task helps to improve the overall quality of experience for users. Indeed, even if the multiple tasks are different models trained by the same FL operator, the operator will not know which model performs best until the training of all models is complete: we therefore would want to dynamically allocate clients to tasks so as to optimize the performance of the worst-performing task, which is accomplished by optimizing for fair task accuracies, as well as high average accuracy. Existing MMFL works that focus only on average task accuracy cannot also ensure fair performance across tasks. This goal of fairness is widely accepted as an important performance criterion for generic computing jobs [18, 19, 20], but is less well-studied in MMFL. To the best of our knowledge, we are the first to extend the MMFL setting to consider fair training of concurrent FL models.
@@ I-A Research Challenges in Fair MMFL
In considering fairness across multiple different models, we seek to ensure that each model achieves comparable training performance, in terms of their converged accuracies, or the time taken to reach threshold accuracy levels. The first challenge to doing so is that the multiple models can have varying difficulties and thus require different amounts of data and computing resources to achieve the same accuracy. Existing MMFL training algorithms, e.g., random allocation, or the round-robin algorithm proposed in [12], allocate clients to tasks uniformly on expectation, or assume that tasks have similar or the same difficulty, e.g., training multiple copies of the same model [13]. However, while such strategies arguably achieve fairness, or at least equality, with respect to resource (i.e., client) allocation to different models, they may lead to “unfair” outcomes. For example, smart home devices may train simpler temperature prediction models that take low-dimensional timeseries as input data, as well as complex person identification ones that use high-dimensional image input data [3]. Allocating equal resources to each model may then yield poor performance for the more difficult ones. As the easier tasks may need less resources in the form of clients compared to the more difficult tasks, a further optimization of client-task allocation could lead to faster convergence and higher accuracies for the more difficult tasks, while maintaining the performance of the easier tasks.
Much work in computer systems and networks has focused on fairness in allocating resources to different computing jobs or network flows, given their heterogeneous resource requirements [18, 19, 20]. Directly applying these allocation algorithms to MMFL, however, is difficult: here, each “resource” is a client, and the performance (e.g. accuracy) improvement realized by allocating this client to a given FL task is unknown in advance and will change over time. This is because it depends on the current model state and the data distribution across clients, both of which are highly stochastic. We must therefore handle uncertainty and dynamics in how much each client contributes to each task’s training performance.
Finally, purely viewing MMFL fairness as a matter of assigning clients to tasks ignores the fact that clients may not be equally willing to contribute resources to each model. For example, smartphone users may benefit from better next-word prediction models if they text frequently, but may derive limited benefit from a better advertisement recommendation model. Also, users may refuse to compute updates if they judge that associated computing costs outweigh personal benefits. If multiple users prefer one model over another, then the disfavored model will overfit to users willing to train it, again leading to an unfair outcome. In an extreme case, no users may be assigned to train this task at all. Prior work[21] has proposed mechanisms to incentivize clients to help train multiple non-simultaneously trained FL models, but their focus was on optimizing the task take-up rate, instead of ensuring that clients are fairly distributed across models.
@@ I-B Our Contributions
Our goal is to design methods to ensure fair concurrent training of MMFL, addressing the challenges above. To do so, we utilize the popular metric of $\alpha$-fairness, commonly used for fair resource allocation in networking [22] and previously proposed to measure fairness across clients in training a single FL model [23]. We use $\alpha$-fairness and Max-Min fairness (a special case of $\alpha$-fairness when $\alpha\rightarrow\infty$ [22]) to design novel algorithms that both incentivize clients to contribute to FL tasks and adaptively allocate these incentivized clients to multiple tasks during training. Intuitively, these algorithms ensure that client effort is spread across the MMFL tasks in accordance with their difficulty, helping to achieve fairness with respect to task performance without compromising average task performance. After discussing related works in Section II, we make the following contributions:
  • We first suppose that all clients are willing to train all models, and propose FedFairMMFL, an algorithm to fairly allocate clients to tasks (i.e., models) in Section III. Our algorithm dynamically and adaptively allocates clients to tasks according to the current performance levels of all tasks. We design FedFairMMFL to approximately optimize the $\alpha$-fairness of all tasks’ training accuracies.
  • From the perspective of a single model, our client-task allocation algorithm in MMFL can be viewed as a form of client selection. With this perspective, we analyze the convergence of MMFL tasks under FedFairMMFL in Section IV, showing that FedFairMMFL preferentially accelerates convergence of more difficult tasks relative to an equal allocation of clients to tasks (Theorem 4 and Corollary 5) and allows each model’s training to converge (Corollary 6).
@@
  • We finally validate our client incentivization and allocation algorithms on multiple FL datasets. Our results show that our algorithm achieves a higher minimum accuracy across tasks than random allocation, round robin and client-fairness baselines, and a lower variance across task accuracy levels, while maintaining the average task accuracy. Our max-min client incentivization further improves the minimum accuracy in the budget-constrained regime.
@@ II Related Work
We distinguish our MMFL setting from that of multi-task learning (MTL), in which multiple training tasks have the same or a similar structure [24], e.g., a similar underlying lower rank structure, and are jointly learnt. Unlike clustered FL [25, 26], our work is different from combining MTL with FL. Instead we consider multiple distinct learning tasks, and their competition for limited client training resources. Our analysis of MMFL performance is closer to prior convergence analysis of client selection in FL [27, 28, 29]. We adapt and modify prior work on FL convergence to reveal the relationships between model convergence and our definition of fairness across tasks.
Past work on multi-model federated learning also aims to optimize the allocation of clients to model training tasks. To the best of our knowledge, we are the first to explicitly consider fairness across the different tasks. Some algorithms focus on models with similar difficulties, for which fairness is easier to ensure [12], while other works propose reinforcement learning [15], heuristic [14], or asynchronous client selection algorithms [16, 11] to optimize average model performance. Other work optimizes the training hyperparameters [30] and allocation of communication channels to tasks [27].
@@ III Fair Client Task Allocation for Multiple-Model FL
We aim to train multiple models (i.e., tasks) [^1] We use model and task interchangeably in this paper. concurrently, such that there is fair training performance across models, e.g., in terms of the time taken, or converged accuracies. Models may be trained on distinct or the same datasets. This performance is a function of which clients train which model at which training round. In this section, we assume that clients have already committed to training all models, and propose a client-model (i.e., client-task) allocation scheme that enables fair training of the multiple tasks. In Sec. V, we introduce auction mechanisms, which incentivize clients to commit to tasks at a fair rate across tasks, before training begins (Fig. 1).
@@
MMFL framework: Consider $K$ clients collaboratively training $S$ distinct machine learning tasks in MMFL (Fig. 1). We let the set of tasks be $\mathbb{S}$. If $S=1$, our setting reduces to traditional FL. Each client is equipped with local data for training each task. All clients can communicate with a central server that maintains a global model $\textbf{w}_{s}$ for each task $s\in\mathbb{S}$. Let each client $k$’s local loss over its local dataset $\mathcal{D}_{k,s}$ for task $s$ be $F_{k,s}(\textbf{w}_{s})=\sum_{d\in\mathcal{D}_{k,s}}\ell\left(d,\textbf{w}_{s}\right)/\left|\mathcal{D}_{k,s}\right|$, where $\ell$ is a loss function that measures the performance of the model $\textbf{w}_{s}$ on data point $d$, and $\left|\mathcal{D}_{k,s}\right|$ is the size of client $k$’s dataset for task $s$. We then define the global loss for task $s$ as a weighted sum over the local losses:
$f_{s}(\textbf{w}_{s})=\sum_{k=1}^{K}p_{k,s}F_{k,s}(\textbf{w}_{s}),$ (1)
where $p_{k,s}=\frac{\left|\mathcal{D}_{k,s}\right|}{\sum_{k}\left|\mathcal{D}_{k,s}\right|}$ is the fraction of training data for task $s$ that resides at client $k$. Similar to traditional FL, in MMFL, the usual aim is to minimize the sum of losses [16]:
@@
Fairness over tasks: We introduce fairness into this objective by utilizing the celebrated $\alpha$-fair utility function often employed to ensure fair network resource allocation [22, 19]. Viewing the loss of each model as an inverse “utility”, we then define our $\alpha$-fair MMFL objective as
$\min_{\textbf{w}_{1},\textbf{w}_{2},...\textbf{w}_{S}}g\left(\textbf{w}_{1},\textbf{w}_{2},...\textbf{w}_{S}\right)\equiv\sum_{s=1}^{S}f_{s}^{\alpha}(\textbf{w}_{s}),$ (3)
@@
To solve (3) within the FL framework, we will design a training algorithm given the gradient
$\nabla_{\textbf{w}_{s}}g=\alpha f_{s}^{\alpha-1}(\textbf{w}_{s})\sum_{k=1}^{K}p_{k,s}\nabla F_{k,s}(\textbf{w}_{s}).$ (4)
FedFairMMFL training (Algorithm 1): Our MMFL training begins with the initialization of the global model weights $\textbf{w}_{s}$ for all tasks, along with an initial random client-task allocation (line 1). The algorithm then enters a loop over global rounds (lines 2-16). During each global round $t$, $m=C\times K$ clients are active (line 3). Here $C$ is the rate at which clients are active. Clients can be inactive due to reasons such as low or unreliable connectivity, devices sleeping to conserve battery, or the lack of incentives on the part of clients. For all active clients, task allocation happens. We let $Sel_{s}(t)$ denote the set of clients allocated to task $s$ in round $t$. Following prior work on MMFL [12, 13, 15], we allocate each client to at most one task during each training round, i.e., the $Sel_{s}(t)$ are disjoint across tasks $s$. For example, mobile clients or other resource constrained devices may be performing other background tasks, e.g., running multiple applications, and thus may not have the computing power or battery to complete training rounds of multiple tasks.
Intuitively, uniform client-task allocation methods, like random or round robin allocation [13], will train simpler tasks faster, since they allocate an equal number of clients to each task, on expectation. We would instead like to ensure tasks are trained at more similar rates (i.e., achieve more fair performance) with a difficulty-aware client-task allocation algorithm, in which more users are allocated to tasks that currently have a higher loss compared to other tasks. We can do so by inspecting the gradient of our MMFL objective in (4). Since resource-constrained clients can only compute gradients for one task per training round, we cannot exactly compute these gradients in (4). Instead, we will assign each client to contribute to the gradient $\nabla F_{k,s}(\textbf{w}_{s})$ for each task $s$ at a probability in proportion to its weight for that task, i.e., proportional to $\alpha f_{s}(\textbf{w}_{s})^{\alpha-1}$. We therefore allocate a task to each active client according to the probability vector p (lines 4-5):
@@
This allocation fulfills our intuition: tasks currently with higher losses $f_{s}$ have a higher probability of being allocated clients. For example, when $\alpha\rightarrow\infty$, all clients will be allocated to the worst performing task, and when $\alpha=1$, allocation will be uniform. Our algorithm does not assume that the task with larger model is the more difficult task, rather it takes the current performance (loss) level $f_{s}(\textbf{w}_{s})$ as a signal on which task needs more clients currently. In Section IV, we formalize this intuition by deriving a convergence bound for each task using FedFairMMFL. This selection scheme is unbiased across clients: for each specific task, every client has the same probability of being selected.
Normalization: Different models’ loss values may belong to different ranges, making them difficult to compare. Therefore, in practice, we ensure comparable values by instead using the accuracy percentages, for all tasks, as an estimate of $f_{s}(\textbf{w}_{s}(t))$, to obtain Eq. (5).
@@
Given the clients allocated to each task $s$ at time $t$, $Sel_{s}(t)$, the server then sends the latest global model weights $\textbf{w}_{s}(t)$ to task $s$’s allocated clients (lines 7). Each client then trains its task on its corresponding local dataset by running $\tau$ SGD steps on the local loss $F_{k,s}$ of its assigned task $s$, as in traditional FL [1] (line 8-9). Let $\textbf{w}_{k,s}(t)$ denote the updated local model for active client $k$ on its allocated task $s$. Each active client then uploads $\textbf{w}_{k,s}(t)$ to the server (line 10), which aggregates the model parameters to synchronize the global model. This aggregation uses weights
$p_{k,{\color[rgb]{0,0,0}Sel_{s}}}=p_{k,s}/\sum_{k^{\prime}\in{\color[rgb]{0,0,0}Sel_{s}}}p_{k^{\prime},s},$ (6)
which indicate the fraction of the training data across selected clients that resides at client $k$, for each active client $k$ (line 13). $Sel_{s}$ is the set of clients training task $s$ in this training round. Hence, $p_{k,{\color[rgb]{0,0,0}Sel_{s}}}$ is a function of $p_{k,s}$. The server then obtains the global loss for each task $f_{s}$, to guide allocation in the next training round (line 14). The next training round then begins.
Reducing overheads: To reduce communication overhead, in FedFairMMFL, we can instead let the central server evaluate the global loss $f_{s}(\textbf{w}_{s})$ on its test set, for client task allocation. This reduces communication and computation (inference) overhead each global round.
Distribution of active clients: For simplicity, we assume that clients have the same active rate. Our algorithm can also be applied for the setting where clients have different active rate from each other and over time: given the active clients each global round, our algorithm will allocate each active client to a task, according to the probability vector p (Eq. (5)).
In Table I, we present the various parameters of our system model.
TABLE I: Definition of Notations
Notations Definitions
$\mathbb{S}=\{1,2,...,s,...,\}$ The set of all FL tasks.
$K$ The total number of clients.
$\mathbf{w}_{s}$ Global model for task $s$.
$\mathcal{D}_{k,s}$ Client $k$’s local dataset for task $s$.
$F_{k,s}(\textbf{w})$ Client $k$’s local loss for task $s$, given model w.
$\ell\left(d,\textbf{w}_{s}\right)$ The loss of model $\textbf{w}_{s}$ on datapoint $d$.
$f_{s}(\textbf{w}_{s})$ Global loss for task $s$.
$p_{k,s}$ Fraction of task $s$’s data residing at client $k$.
$f_{s}^{\alpha}(\textbf{w}_{s})$ $\alpha$-fair global loss function for task $s$.
$g(\textbf{w}_{1},\textbf{w}_{2},...)$ Sum of the $\alpha$-fair global loss function, across tasks.
$\mathbf{p}$ Client selection probability vector.
$Sel_{s}(t)$ Subset of clients allocated to task $s$ at round $t$.
$\textbf{w}_{k,s}(t)$ Local model, client $k$, task $s$, at time $t$.
$p_{k,Sel_{s}}$
Proportion of the data which client $k$ has, amongst
the selected clients in $Sel_{s}$.
$a_{k,s}(t)$ Variable indicating if client $k$ trains task $s$.
$B^{s}_{Sel_{s}}(\alpha)$
Probability that a specific set of clients $Sel_{s}$ is
selected for task $s$ by our algorithm. See Eq. (16).
$\textbf{w}_{s}^{*}$ Global optimal.
$\textbf{w}_{k,s}^{*}$ Client $k$’s optimal.
$\Gamma_{s}$
The gap between the loss evaluated at the global
optimum $\mathbf{w}_{s}^{*}$, and the loss evaluated at $\mathbf{w}^{*}_{k,s}$,
for task $s$. See Eq. (17).
$\rho^{\alpha}(B^{s}_{Sel_{s}},\textbf{w}^{\prime})$
Selection skew, the ratio of the expected ’local-
global objective gap’ of our algorithm’s selected
clients, over the ’local-global objective gap’ of all
$K$ clients. See Eq. (18).
$\overline{\rho}^{\alpha}$, $\underline{\rho}^{\alpha}$ Upper and lower bounds on the selection skew.
$\sigma^{2}$ Bound on the variance of stochastic gradient.
$(\mathbf{w}^{*}_{1_{\alpha}},...\mathbf{w}^{*}_{S_{\alpha}})$
The global optimal solution when parameter $\alpha$ is
used.
@@
2:for $Global~{}epoch~{}t=1,2,\ldots T$ do
3:    $m=C\times K$ clients are active.
@@
7:         Server sends the latest model weights of task $s$ ($\textbf{w}_{s}$).
8:         Client performs local training for $\tau$ local rounds for
9:         allocated task, optimizing $\textbf{w}_{k,s}$.
10:         Client uploads updated $\textbf{w}_{k,s}$ to the server.
11:    end for
12:    for all tasks $s\in\mathbb{S}$ do
13:         Server aggregates $\textbf{w}_{s}\leftarrow\sum_{k\in Sel_{s}}p_{k,Sel_{s}}\textbf{w}_{k,s}$.
14:         Obtain global accuracy $f_{s}$ for task $s$.